find the value of x so that the points lie on a line with a given slope:
(x,2) and (-5,8), m = -1

Answer by mananth(12270)

(x,2) and (-5,8), m = -1
..
slope = y2-y1/x2-x1
-1=(8-2)/(-5-x)
-1=6/-5-x
-1(-5-x)=6
5+x=6
-5 -5
x=1
...
m.ananth@hotmail.ca